### v3

# MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling

Abstract.
Despite recent advances, long-sequence video generation frameworks still suffer from significant limitations: poor assistive capability, suboptimal visual quality, and limited expressiveness. To mitigate these limitations, we propose MAViS, an end-to-end multi-agent collaborative framework for long-sequence video storytelling. MAViS orchestrates specialized agents across multiple stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.In each stage, agents operate under the 3E Principle—Explore, Examine, and Enhance—to ensure the completeness of intermediate outputs. Considering the capability limitations of current generative models, we propose the Script Writing Guidelines to optimize compatibility between scripts and generative tools. Experimental results demonstrate that MAViS achieves state-of-the-art performance in assistive capability, visual quality, and video expressiveness. Its modular framework further enables scalability with diverse generative models and tools. With just a brief user prompt, MAViS is capable of producing high-quality, expressive long-sequence video storytelling, significantly improving the assistive capability of enriching inspirations and creativity for users. To the best of our knowledge, MAViS is the only framework that provides multimodal design output – videos with narratives and background music.
$\dagger$ Corresponding author and project lead
1. Introduction
In recent years, video generative models (OpenAI, 2024; MiniMax, 2024; GenmoTeam, 2024) have made substantial strides in text-to-video (T2V) and image-to-video (I2V) generation. While these methods demonstrate promising capabilities for shot-level creation, the generation of long videos remains a formidable challenge. Current scale of compute and data is insufficient to train a single end-to-end model capable of producing high-quality, minute-long videos.
Table 1. Properties of MAViS vs Other Methods: We summarize seven metrics for long-sequence video storytelling works.”Agentic”: whether the system employs a multi-agent architecture; ”Script Writing”: ability to automatically generate scripts from user prompt; ”Shot Designing.”: support for structured shot designing and controllable narrative flow; ”Dynamic”: whether the generated video exhibits dynamic features; ”ID Consistency”: consistency in character identity across multiple shots; ”Audio”: whether the output includes synchronized audio; ”Automated”: whether the pipeline can operate without the presence of user input in script writing, script planning, or model fine-tuning.
Methods Agentic Script Writing Shot Designing Dynamic ID Consistency Audio Automated
Mora (Yuan et al., 2024) $\times$ $\times$ $\times$ $\times$
AesopAgent (Wang et al., 2024c) $\times$ $\times$ $\times$
MovieDreamer (Zhao et al., 2024) $\times$ $\times$ $\times$ $\times$ $\times$
DreamFactory (Xie et al., 2024) $\times$ $\times$ $\times$
VGoT (Zheng et al., 2025) $\times$ $\times$ $\times$
LCT (Guo et al., 2025) $\times$ $\times$ $\times$ $\times$ $\times$
MovieAgent (Wu et al., 2025) $\times$ $\times$
MAViS (ours)
To overcome this limitation, prior works have explored extending existing video clips (KlingAI, 2024; Henschel et al., 2024; Ge et al., 2022), or composing sequences of keyframes (Liu et al., 2024; Yang et al., 2024a; Rahman et al., 2023; Wang et al., 2024c), or integrating storytelling and video generation (Yin et al., 2023b; Zheng et al., 2025). However, these approaches often suffer from repetitive motion, lack of narrative cohesion, or result in static image series with limited visual expressions. To better align narrative planning with video animation, MovieDreamer (Zhao et al., 2024) adopts an auto-regressive approach to generate key frames from input scripts, which are then expanded into videos via I2V models, while Dreamfactory (Xie et al., 2024) and MovieAgent (Wu et al., 2025) propose a hierarchical script–scene–shot pipeline. However, as shown in Table 1, these methods require substantial manual effort from users for script writing and training LoRA models to maintain character identity consistency, which significantly limits their assistive capability and hinders large-scale deployment.
In summary, current long-sequence video storytelling methods face three major limitations: 1) poor assistive capability, not supporting script writing, shot designing, and LoRA training to maintain character ID consistency; 2) suboptimal visual quality, manifested in visual distortions, unnatural actions, and inaccurate object proportions; and 3) limited expressiveness, often resulting in repetitive actions and poorly structured shot compositions.
To address this, we introduce MAViS, a multi-agent framework for end-to-end long-sequence video generation. MAViS encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation. Given a brief user prompt, the script writing stage first generates a structured script with character definitions and shot outlines. The shot designing stage defines a series of shot-level elements to guide subsequent content generation, while character modeling ensures identity consistency via LoRA-trained appearances. During the keyframe generation stage, T2I models are employed to create the initial static frames for each shot, i.e. the keyframes, which are then expanded into full video sequences in the video animation stage through I2V models, maintaining visual coherence and narrative fluency. Finally, voice-overs and background music are automatically added to each shot during the voice generation stage, resulting in a cohesive long-sequence video.
Crucially, all stages operate under the 3E Principle: - Explore: Initial generation of candidate outputs; - Examine: Reviewing the quality and completeness of the output; - Enhance: Iterative refinement and optimization. After multiple iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the end-to-end pipeline.
To mitigate the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, we propose the Script Writing Guidelines aimed at enhancing the compatibility between scripts and generative models, ultimately improving the expressiveness of the video storytelling. With just a brief user prompt, MAViS efficiently generates high-quality and complete long-sequence video storytelling, significantly improving the assistive capability of enriching inspirations and creativity for users.Furthermore, its modular architecture ensures scalability with various generative tools and models.
Our contributions are summarized in three thrusts:
$\bullet$ We propose MAViS, an end-to-end multi-agent framework that encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation — forming a complete and scalable workflow for long video storytelling.
$\bullet$ We introduce the 3E PrincipleExplore, Examine, Enhance—to continuously optimize the output of each stage. This iterative design ensures the completeness of the output at each stage.
$\bullet$ We propose the Script Writing Guidelines for long-sequence video storytelling to enhance the script-tool compatibility, ultimately improving the expressiveness of the video storytelling.
$\bullet$ Experiments demonstrate that MAViS achieves state-of-the-art performance and win the most user preferences in terms of expressiveness and visual quality. With just a brief user prompt, MAViS efficiently generates high-quality and complete long-sequence video storytelling. To the best of our knowledge, MAViS is the only framework that provides multimodal design output.
2. Related Works
2.1. Image Generation
Recent advances in image generation (Li et al., 2019; Xu et al., 2018; Zhang et al., 2017; Ramesh et al., 2021) have been propelled by deep learning techniques.To fulfill the task of visual storytelling, some works (Liu et al., 2024; Yang et al., 2024a; Rahman et al., 2023) generate image sequences from text and image-caption pairs. To ensure subject consistency (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) adopts subject-specific fine-tuning, while GPT-Image-1 (OpenAI, 2025) introduces context-aware generation for identity preservation. In this work, MAViS leverages off-the-shelf models to generate keyframes for each shot, which serves as the basis for subsequent video animation.
Figure 2. Illustration of MAViS framework. See additional presentation is in the supplementary video.
2.2. Video Generation
Long video generation (He et al., 2022; Ge et al., 2022; Wang et al., 2024e) aims to produce longer videos than previous text-to-video (T2V) (ModelScopeTeam, 2023; Wang et al., 2024b) and image-to-video (I2V) (Girdhar et al., 2023; Xing et al., 2025; Zhang et al., 2023) models. Auto-regressive approaches (Weng et al., 2024; Henschel et al., 2024) extend videos by conditioning on earlier frames, yet often result in repetitive motions with limited narrative development. Attempts to directly integrate storytelling and video generation (Yin et al., 2023b; Zheng et al., 2025) suffer from poor visual quality and unstructured video composition, leading to subpar viewer experiences. Meanwhile, commercial models have achieved notable success in both T2V (GenmoTeam, 2024; Kong et al., 2024; OpenAI, 2024) and I2V generation (Google, 2024; Luma, 2024; MiniMax, 2024; Ang and Baole, 2025), and some  (KlingAI, 2024; Runway, 2024a, b) support temporal extension. However, the duration of generated videos remains constrained. In this work, MAViS uses cutting-edge video generation models to synthesize high-quality video clips and assembles them into long-sequence video storytelling.
2.3. AI Agents
AI agents, empowered by large language models (Brown et al., 2020; Touvron et al., 2023a, b), multimodal architectures, and reinforcement learning, have shown broad applicability in reasoning, planning, web interaction, and visual generation. General-purpose LLMs underpin agent frameworks (Yang et al., 2023b; Yin et al., 2023a; Li et al., 2023; Hong et al., 2023; Park et al., 2023), while multimodal models (Achiam et al., 2023; Anil et al., 2023) extend agent capabilities across modalities (Yang et al., 2023a; Zhang and Zhang, 2024; Zheng et al., 2024; Wang et al., 2024a; Liu et al., 2025; Wang et al., 2024d; Hang et al., 2024; Liao et al., 2025). In video generation, systems (Yang et al., 2024b; Yuan et al., 2024; Wang et al., 2024c) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines like Dreamfactory (Xie et al., 2024) and MovieAgent (Wu et al., 2025) introduce script-to-shot workflows but still rely on manual scripting and LoRA tuning. In contrast, our MAViS leverages multi-agent collaboration to enable long-sequence video storytelling.
3. MAViS
We propose MAViS, a multi-agent end-to-end framework for long-sequence video storytelling that consists of various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.
Given a user prompt $P$, MAViS aims to generate a long-sequence video storytelling that comprises $N$ shots:
(1) $\mathcal{F}:P\rightarrow\hat{\mathcal{V}},\quad\hat{\mathcal{V}}=\{\{{V}_{j},\ {A}_{j}\}\mid j=1,2,\dots,N\},$
where $\hat{V}_{j}$ is the video clip and $\hat{A}_{j}$ is the audio track. To ensure the completeness and reliability of the generation pipeline, agents in each stage follow a unified 3E Principle: Explore, Examine, and Enhance. Section 3.1 details the 3E principle, and Section 3.2 outlines the Script Writing Guidelines, followed by Section 3.3 to 3.6 introducing the agent design and working procedures of each stage.
3.1. 3E Principle
Existing frameworks for long-sequence video storytelling adopt one-shot generation strategies. However, such approaches often prove inadequate in practice—scripts may not align with user intent, images or videos can exhibit distortions or unnatural actions, and audio tracks may be misaligned in duration. These issues highlight the necessity of adopting a multi-round generation–modification loop, rather than relying on one-shot generation. To address this, we introduce the 3E Principle—Explore, Examine, Enhance—to guide intra-stage agents’ collaboration and progressively improve generation quality across iterations:
Explore: This stage represents the first attempt to generate content based on the current request and specification:
(2) $r_{1}=G(p_{1};g),$
where $p_{1}$ denotes the initial generation request, $g$ is the generation guidelines, such as guides of script-tool compatibility and output completeness. $G$ is the generator, and $r_{1}$ is the resulting output.
Examine: The generated output is then examined to determine whether it satisfies the specification:
(3) $f_{1}=E(r_{1},g),$
where $E$ is the reviewer, and $f_{1}$ is the examination feedback of result $r_{1}$ based on the guidelines $g$.
Enhance: Based on the feedback and guidelines, the generation request is refined to improve quality in the next iteration:
(4) $p_{i+1}=R(p_{i};f_{i},g),$
where $R$ denotes the refiner, and $p_{i+1}$ is the updated request for iteration $i+1$. This new request serves as the input to the next round of Explore-Examine-Enhance, and the process continues iteratively until the output fully satisfies the specification.
After multiple Explore-Examine-Enhance iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the pipeline. Given the limitations of individual agents, it may need to deploy multiple reviewers in parallel to ensure comprehensive assessment. Additionally, generating multiple candidates per iteration and enhancing the highest-scoring one can effectively reduce the total number of required iterations. Moreover, depending on the design, the generator and refiner may be implemented as a unified component.
3.2. Script Writing Guidelines
Given the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, certain scripts remain infeasible for video generation. To address this gap and improve the compatibility between scripts and generative models, we propose the Script Writing Guidelines for long-sequence storytelling. The Script Writing Guidelines comprises three components: the Structure Guide, which focuses on how scenes and shots are segmented and arranged; the Content Guide, which defines how much content a single shot should include; and the Style Guide, which ensures user-centric alignment, coherent pacing, logical flow, and structural completeness. Together, these guidelines constrain agent-based script writing and improve the expressiveness of the ultimate video storytelling.
Structure Guide
Current video generation models have limited contextual modeling capacity; as a result, adjacent shots set within the same spatial environment often suffer from inconsistency in background elements. Therefore, scriptwriters should avoid placing successive shots in the same background and should minimize transitions between tightly connected spatial locations (e.g., from ”outside the door” to ”inside the room,” or from one room to another). Moreover, to reduce the jarring effect of spatial jumps, it is recommended not to place the same character in consecutive shots unless narrative necessity dictates it.
To mitigate these issues while preserving narrative flow, transitional shots can be inserted. These may include close-ups of different characters, partial views of a character (e.g., hands, back), detailed shots of the environment, or important props. Such insertions help avoid background duplication while reinforcing the spatial or character continuity. Additionally, enhanced descriptions of both background and character appearance are encouraged to improve model alignment with visual and narrative consistency.
Content Guide
Current video generation models struggle to synthesize multiple sequential actions within a single shot (e.g., running followed by walking and then dancing). Therefore, it is advisable to limit each shot to a single, simple action. Furthermore, generative models often fail to produce fine-grained visual elements accurately, such as textual content on smartphone screens or small patterns on objects. Scripts should avoid specifying such details.
Style Guide
Within the Structure Guide and the Content Guide, a script may risk becoming fragmented or visually monotonous if not properly orchestrated. The style guide thus ensures harmonious pacing, visual diversity, and narrative logic. Moreover, the script’s tone and theme must align with the user’s request, and different genres require distinct narrative pacing—emotional and family-oriented scripts benefit from a slow, layered progression, while action-oriented scripts demand rapid and dynamic developments. Additionally, structural completeness is essential: every script must contain a title, character definitions, and a shot-by-shot script. Each shot should specify the character (or ”None” if absent), shot content, and accompanying subtitle or voice-over. Subtitles must align semantically with the shot content.
3.3. Script Writing Stage
Upon receiving a brief user prompt, four agents—the Scriptwriter, the Structure Reviewer, the Content Reviewer, and the Style Reviewer—work collaboratively in the script writing stage to produce a structured script.
The Scriptwriter $G_{\textrm{scr}}$ first generate a draft script $S_{1}$ based on the user prompt $P$ and the Script Writing Guidelines $g_{\textrm{scr}}$:
(5) $S_{1}=G_{\textrm{scr}}(P,g_{\textrm{scr}}).$
Following the 3E principle, at each iteration, the Structure Reviewer $E_{\textrm{str}}$, Content Reviewer $E_{\textrm{con}}$, and Style Reviewer $E_{\textrm{sty}}$ examine the script $S_{i}$ against the Structure Guide $g_{\textrm{str}}$, Content Guide $g_{\textrm{con}}$ and Style Guide $g_{\textrm{sty}}$ respectively, and provide feedback:
(6) $f_{\mathrm{scr}}^{i}=\left[E_{\textrm{str}}(S_{i},g_{\textrm{str}}),\ E_{\textrm{con}}(S_{i},g_{\textrm{con}}),\ E_{\textrm{sty}}(S_{i},g_{\textrm{sty}})\right],$
where $g_{\textrm{scr}}=[g_{\textrm{str}},\ g_{\textrm{con}},\ g_{\textrm{sty}}]$. The Scriptwriter integrates these feedback to improve the script iteratively:
(7) $S_{i+1}=G_{\textrm{scr}}(f_{\mathrm{scr}}^{i};\ S_{i},g_{\textrm{scr}})$
The Explore-Examine-Enhance loop continues until all reviewers reach a consensus that the script fully satisfies the guidelines after $n$ iterations, and the Scriptwriter outputs the finalized version, represented as $S=C_{\textrm{scr}}(S_{n})$.
3.4. Shot Designing Stage
In the shot-designing stage, the Shot Designer—responsible for generating detailed shot designs—and the Shot Reviewer—tasked with assessing the completeness and adherence of shot designs to predefined guidelines—play pivotal roles in transforming the high-level script into detailed shot elements suitable for downstream image and video generation. In detail, each shot is enriched with seven essential elements to enable precise visual control:
$\bullet$ Background: Specifies the spatial and environmental context of the shot, including location, time of day, season, weather conditions, and foreground-background composition.
$\bullet$ Character Pose: Describes the initial static state of characters, including posture, expression, and position when the shot begins.
$\bullet$ Character Action: Captures dynamic interactions or movements of characters throughout the shot.
$\bullet$ Prop Description: Defines non-character objects, including their form, texture, behavior, and placement.
$\bullet$ Camera Position: Specifies the camera’s initial angle, orientation, and distance from the scene.
$\bullet$ Camera Movement: Describes the movement mode and direction of the camera during the shot.
$\bullet$ Lighting Design: Defines lighting characteristics such as type, tone, brightness, and color temperature to convey atmosphere.
Similarly, following the 3E Principle, the Shot Designer generates an initial shot design from the script:
(8) $\hat{S}_{1}=G_{\mathrm{shot}}(S),$
where the shot designing guide provides essential constraints to ensure completeness and stylistic alignment with the original script.
In each iteration $i$, the Shot Reviewer examines the completeness and appropriateness of the current design $\hat{S}_{i}$ and provides feedback:
(9) $f_{\mathrm{shot}}^{i}=E_{\mathrm{shot}}(\hat{S}_{i}).$
The Shot Designer then uses this feedback to enhance the design:
(10) $\hat{S}_{i+1}=G_{\mathrm{shot}}(f_{\mathrm{shot}}^{i};\ \hat{S}_{i}).$
The process repeats until the Shot Reviewer deems the result satisfactory. The final shot design is denoted as $\hat{S}=\hat{S}_{n}$.
3.5. Character Modeling Stage
To ensure the identity consistency of characters across shots, in the character modeling stage, agents construct visual representations for each character and train a LoRA model. This process incorporates the Character Prompt Generator and Caption Generator, supported by reviewer agents for both image generation (T2I Evaluator and T2I Refiner) and video animation (I2V Evaluator and I2V Refiner).
The process begins with the Character Prompt Generator, which generates both text-to-image (T2I) and image-to-video (I2V) prompts for each character. T2I prompts aim to generate a frontal and naturally posed character image, while I2V prompts are used—together with the T2I images—to synthesize character-centric video clips that capture the subject from multiple angles. After both T2I and I2V processes, multi-view character images are sampled from the generated character video. The Caption Generator then produces structured captions for each sampled frame. These images and captions are subsequently used to train LoRA models (if needed).
Both the T2I and I2V processes follow the 3E Principle, but differ from script writing and shot designing in that each iteration produces multiple samples from the generative model pools. Specifically, in the $i$-th iteration, a set of candidate images is generated by the pool of T2I models $\mathcal{M}_{I}$ using the current T2I prompt $p_{I}^{i}$:
(11) $\mathcal{I}_{i}=\{G_{I}^{k}(p_{I}^{i}),\ G_{I}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{I}|\}.$
T2I Evaluator $E_{I}$ scores candidates according to the image evaluation guide $g_{I}$, selects the best one $I_{i}$, and provides evaluation result $f^{i}_{I}$:
(12) $I_{i},f^{i}_{I}]=E_{I}(\mathcal{I}_{i},g_{I})$
The T2I Refiner $R_{I}$ then updates the prompt accordingly:
(13) $p_{I}^{i+1}=R_{I}(f^{i}_{I},p_{I}^{i})$
After $n$ iterations, the final image $I=I_{n}$ initiates the I2V process.
The I2V process mirrors the T2I process. First, a set of candidate videos is generated by the I2V model pool $\mathcal{M}{I}$:
(14) $\mathcal{V}_{i}=\{G_{V}^{k}(I,p_{V}^{i}),\ G_{V}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{V}|\}.$
The I2V Evaluator $E_{V}$ evaluates these candidates according to the video evaluation guide $g_{I}$, selects the best one $V_{i}$, and provides evaluation result $f^{i}_{V}$. The I2V Refiner $R_{V}$ updates the I2V prompt and produce the final video $V_{n}$ after $n$ iterations:
(15) $V_{i},f^{i}_{V}]=E_{V}(\mathcal{V}_{i},g_{V}),\quad p_{V}^{i+1}=R_{V}(f^{i}_{V},p_{I}^{i}),\quad V=V_{n}.$
Table 2. Evaluation of Automatic Metrics for Keyframe Generation and Video Animation. ”T. Flick.”, ”M. Smooth.”, ”Sub. Cons.”, ”Bg. Cons.”, ”Aesthetic”, and ”I. Quality” refer to ”Temporal Flickering”, ”Motion Smoothness”, ”Subject Consistency”, ”Background Consistency”, ‘Aesthetic Quality’, and ”Imaging Quality” from metrics in VBench (Huang et al., 2024), respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
Method Keyframe Generation Video Animation
CLIP $\uparrow$ Inception $\uparrow$ T. Flick. $\uparrow$ M. Smooth. $\uparrow$ Sub. Cons. $\uparrow$ Bg. Cons. $\uparrow$ Aesthetic $\uparrow$ I. Quality $\uparrow$
VGoT (Zheng et al., 2025) 22.37 8.53 99.00 99.31 98.94 98.74 80.11 64.59
Mora (Yuan et al., 2024) 33.98 12.68 97.32 99.23 95.23 94.88 64.36 71.48
MovieAgent (Wu et al., 2025) 31.71 10.72 97.07 99.20 93.29 94.51 63.39 71.26
MAViS (ours) 34.22 12.81 99.09 99.53 95.72 96.12 63.17 72.91
Table 3. User Study on the Performance of Long-Sequence Video Storytelling (Voting Results). ”Narrative”, ”Visual”, ”User Align.”, ”Sub. Cons.”, ”Sub. Natural.”, ”Bg. Cons.”, and ”Bg. Real.” refer to ”Narrative Expressiveness”, ”Visual Quality”, ”User Prompt Alignment”, ”Subject Consistency”, ”Character Naturalness”, ”Background Consistency”, and ”Background Realism”, respectively. $\uparrow$ indicates a higher value is more desired. Bold indicates the best results.
Method Narrative $\uparrow$ Visual $\uparrow$ User Align. $\uparrow$ Sub. Cons. $\uparrow$ Sub. Natural. $\uparrow$ Bg. Cons. $\uparrow$ Bg. Real. $\uparrow$
VGoT (Zheng et al., 2025) 3.39 3.39 6.79 5.56 3.07 4.14 3.24
Mora (Yuan et al., 2024) 15.18 16.96 13.57 15.77 13.00 13.69 15.29
MovieAgent (Wu et al., 2025) 9.46 11.79 11.96 12.37 12.27 12.07 10.97
MAViS (ours) 71.96 67.86 67.68 66.31 71.66 70.09 70.50
The difference is that the I2V process incorporates additional evaluation complexity. The image evaluation guide assesses candidates based on three primary axes: 1) Visual Quality: clarity, detail resolution, and aesthetic harmony; 2) Naturalness: identifying distortions, anatomical inaccuracies, or proportion inconsistencies; and 3) Prompt Consistency: if the image content aligns with the input prompt. In contrast, the video evaluation guide extends these criteria by introducing two additional dimensions: 1) Subject Consistency: temporal stability in facial features, attire, hairstyle, and skin tone across frames; and 2) Dynamics: the richness, fluidity, and pacing of character movement throughout the clip. Together, the image and video evaluation guides work in concert to enforce the completeness of the final generation.
3.6. Generation Stages
The generation stages include: 1) the keyframe generation stage, which comprises the T2I Prompt Generator, T2I Evaluator, and T2I Refiner, responsible for generating the initial reference image for each shot (i.e., the keyframes); 2) the video animation stage, which includes the I2V Prompt Generator, I2V Evaluator, and I2V Refiner, and completes the video shot based on the keyframes; and 3) the audio generation stage, which consists of the Voice Designer, Voice Reviewer, and Subtitle Refiner, synthesizes voice-overs and background music aligned with the emotional rhythm of each shot.
First, in the keyframe generation stage, the T2I Prompt Generator uses shot-level elements—Character, Background, Character Pose, Prop Description, Camera Position, and Lighting Design—to construct T2I prompts, which are refined via the same T2I process in Section 3.5, utilizing the T2I model pool and the trained LoRA model, resulting in a set of keyframes $\{I_{j}\mid j=1,2,\dots,N\}$.
Second, in the video animation stage, the I2V Prompt Generator formulates I2V prompts using Character, Background, Character Action, Prop Description, Camera Movement, and Lighting Design. Again, following the same I2V process introduced in Section 3.5, these prompts are used to generate video clips $\{V_{j}\mid j=1,2,\dots,N\}$.
After video animation, voice and music planning are initiated in the audio generation stage based on the 3E principle. The Voice Designer $G_{A}$ selects background music for the entire script and configures voice IDs and emotional tones for each individual shot. The voice design is then reviewed by the Voice Reviewer $E_{A}$, who checks for compliance (e.g., music availability and voice–emotion appropriateness) and provides iterative feedback $f_{A}^{i}$ to assist in refinement:
(16) $S_{A}^{1}=G_{A}(S),\quad f_{A}^{i}=E_{A}(S_{A}^{i}),\quad S_{A}^{i+1}=G_{A}(S,f_{A}^{i}),\quad S_{A}=S_{A}^{n}.$
Next, the T2A model $G_{A}$ generates speech audio for each shot. The Subtitle Refiner $R_{A}$ shortens the voice-over $p_{A}$ to fit temporal constraints, ensuring compatibility with video animation:
(17) $A^{i}_{j}=G_{A}(p_{A,j}^{i}),\quad p_{A,j}^{i+1}=R_{A}(p_{A,j}^{i};A^{i}_{j},V_{j}),\quad A_{j}=A^{n}_{j}.$
Finally, audio tracks are synchronized with video clips to produce the final long-sequence video telling with synchronized visuals and sound: $\hat{\mathcal{V}}=\{\{V_{j},\ A_{j}\}\mid j=1,2,\dots,N\}$.
4. Experiment
To evaluate the performance of MAViS, we compared it with three open-sourced long video generation frameworks and conducted fine-grained ablation studies. Further experimental details and results are provided in the supplementary material.
Implementation Details
All agents in our framework are implemented using AutoGen (Wu et al., 2024) and powered by GPT-4o (Achiam et al., 2023). The T2I model pool comprises FLUX.1 (BlackForestLabs, 2023), Stable Diffusion 3.5 (Rombach et al., 2022), and GPT-Image (OpenAI, 2025), while the I2V model pool includes Veo2 (Google, 2024), Gen-3 (Runway, 2024a), Gen-4 (Runway, 2024b), LUMA (Luma, 2024), HunyuanVideo-I2V (Kong et al., 2024), and Wan2.1 (Ang and Baole, 2025). We use MUDI (Jang et al., 2024) for LoRA training and Hailuo (MiniMax, 2024) for the T2A generation. The maximum number of 3E iterations is set to 4 for script writing, shot designing, and voice designing; 2 for the T2I process; 1 for the I2V process; and 5 for subtitle refining. As the first work targeting end-to-end long-sequence video storytelling, no suitable benchmark dataset exists. Therefore, we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation. All experiments were conducted on one NVIDIA V100 GPUs.
Metrics
We employ CLIP and Inception Score to evaluate prompt consistency and generation quality in the keyframe generation stage, and use VBench (Huang et al., 2024) to assess video animation performance. Additionally, a user study is conducted to assess real user preferences and validate the practical effectiveness of frameworks. For all evaluation metrics, higher values indicate better performance.
Baselines
We compare MAViS with three long video generation baselines: VGoT, Mora, and MovieAgent. Among them, Mora and MovieAgent adopt a modular design and are equipped with the same image and video generation models as MAViS, and are also powered by GPT-4o. Since Mora and MovieAgent do not support script generation, we provide them with scripts produced by our script writing stge. For MovieAgent, which requires manual training of LoRA models, we generate character images based on the provided scripts and train LoRA models for each script.
Table 4. Ablation study on the Structure Reviewer $E_{\mathrm{str}}$, the Content Reviewer $E_{\mathrm{con}}$, and the Style Reviewer $E_{\mathrm{sty}}$. $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower than the others.
Ablation Narrative $\uparrow$ Visual $\uparrow$ User Align. $\uparrow$
w/o $E_{\mathrm{str}}$ 13.67 25.67 26.67
w/o $E_{\mathrm{con}}$ 32.33 18.67 26.67
w/o $E_{\mathrm{sty}}$ 20.00 26.67 20.00
MAViS (ours) 34.00 26.00 26.67
4.1. Performance Comparisons
We evaluate each method via automatic metrics and user study. It’s worth noting that MAViS is the only framework that provides multimodal output – videos with narratives and background music.
4.1.1. Automatic Metrics
Table 2 presents the shot-level evaluation results for both keyframe generation and video animation. MAViS achieves the highest scores in CLIP (4.22) and Inception (12.81) scores, attributed to the enhanced T2I quality enabled by the 3E Principle in the T2I process. In video animation, MAViS slightly lags behind VGoT in subject consistency and background consistency but outperforms all baselines in temporal flickering, motion smoothness, and image quality. This is partly because VGoT produces videos with limited dynamics and highly saturated visuals, which improves intra-shot consistency and aesthetic quality but at the cost of realism. Overall, MAViS achieves the best comprehensive performance.
4.1.2. User Study
Table 3 summarizes the voting results from 60 evaluators across multiple aspects of comparison among the baselines. MAViS consistently outperforms all other methods in every evaluation dimension, with an average voting share of 69.44%. This highlights both the effectiveness of our architecture and the critical role of the 3E Principle. Detailed settings of the user study are provided in the supplementary material.
4.2. Ablation Study
We conduct an ablation study to demonstrate the impact of the 3E Principle across different modules of the framework.
4.2.1. Script Writing Guidelines
In the script writing stage, based on the Script Writing Guidelines, the Structure Reviewer focuses on how scenes and shots are segmented and arranged, the Content Reviewer ensures that each shot can fully convey its intended content, and the Style Reviewer oversees global pacing, logical flow, structural coherence, and alignment with user requirements. Table 4 presents the effects of removing each reviewer from the pipeline. As shown, excluding the Structure Reviewer significantly impairs narrative expressiveness, removing the Content Reviewer degrades visual quality, and omitting the Style Reviewer leads to a loss in both narrative alignment and consistency with user intent. Optimal script generation is achieved only when all three reviewers collaborate.
Table 5. Ablation study of the 3E Principle in the T2I and I2V Processes. ”Natural.” refers to the Naturalness score, $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
Ablation CLIP $\uparrow$ Inception $\uparrow$ Natural. $\uparrow$ Sub. Cons. $\uparrow$ Dyn. Deg. $\uparrow$
T2I w/o 3E 31.25 10.54 72.00 - -
w 3E 34.22 12.81 98.00 - -
I2V w/o 3E 32.05 11.22 60.00 3.0 44.58
w 3E 34.53 13.22 92.50 95.72 35.62
Table 6. Ablation study on the Shot Reviewer $E_{\textrm{shot}}$, the Voice Reviewer $E_{A}$, and the Subtitle Refiner $R_{A}$. ”C. Rate” refers to ”Compliance Rate”. $\uparrow$ indicates a higher value is more desired. Gray-shaded cells indicate values that are significantly lower.
Ablation C. Rate $\uparrow$ Ablation C. Rate $\uparrow$ Ablation C. Rate $\uparrow$
w/o $E_{\textrm{shot}}$ 84.5 w/o $E_{A}$ 92.0 w/o $R_{A}$ 56.5
w $E_{\textrm{shot}}$ 100.0 w/o $E_{A}$ 100.0 w/o $R_{A}$ 100.0
4.2.2. 3E Principle in T2I and I2V Processes
Since the evaluators and refiners in both the T2I and I2V processes cannot operate independently, we conduct an ablation study to evaluate the overall impact of the 3E Principle on these modules. The CLIP score, Inception score, Naturalness score (defined as the proportion of generated images or videos that appear physically plausible and naturally rendered), along with Subject Consistency and Dynamic Degree from the VBench metrics, are used to assess key aspects targeted by the T2I and I2V Evaluators—specifically, prompt consistency, visual quality, naturalness, subject consistency, and dynamics. As shown in Table 5, removing the 3E Principle leads to a noticeable drop in performance across both T2I and I2V outputs, particularly in terms of naturalness ($72.0\rightarrow 98.0$ in T2I and $60.0\rightarrow 92.5$ in I2V). This highlights the critical role of the 3E Principle in improving generation quality. An exception is observed in the I2V process: the Dynamic Degree is higher without the 3E Principle ($44.58>35.63$). This may be because high-dynamic videos are more prone to physical inconsistencies and instability, leading the agents to favor lower-dynamic, more stable results when guided by the 3E Principle.
4.2.3. Shot Reviewer, Voice Reviewer, and Subtitle Refiner
As related, the Shot Reviewer assesses the completeness of the shot designs, the Voice Reviewer ensures the music availability and voice-emotion appropriateness of voice design, and the Subtitle Refiner shortens voice-over to match the duration limits of video clips. We use Compliance Rate—defined as the proportion of outputs that fully adhere to their respective guidelines—to evaluate the contributions of these agents. As shown in Table 6, removing these agents leads to a noticeable drop in compliance, particularly in voice-over generation, which in turn negatively impacts downstream content generation. These results underscore the necessity of the 3E Principle in shot designing, voice designing, and the speech audio generation.
5. Conclution
In this work, we presented MAViS, a multi-agent end-to-end framework for long-sequence video storytelling. By orchestrating multiple stages within a unified pipeline, MAViS effectively addresses the key limitations of prior approaches — poor assistive capability, limited visual expressiveness, and incomplete outputs. Central to our framework is the 3E Principle (Explore, Examine, Enhance), which enables iterative refinement across all stages, maximizing output quality. Additionally, our proposed Script Writing Guidelines improve compatibility between narrative structure and generative tools. Experimental results validate that MAViS significantly advances the assistive capability, visual quality, and expressiveness of long-sequence video storytelling.